OpenStudy (anonymous):

An 85m long train begins accelerating at a constant rate from rest. The front of the train passes a railway worker, who is standing 300 meters down the track from where the front of the train started, at a speed of 35m/s. What will the speed of the last car be when it passes the worker?

OpenStudy (roadjester):

we don't know the acceleration?

OpenStudy (anonymous):

Tricky question. Get the acceleration from the increase in kinetic energy, which gives us v^2 = 2 a s, with v = velocity = 35 m/s a= acceleration (F/m) s = distance = 300 m. When the last car passes by, the first one will have gone s = 385 m, and the same equation gives the velocity of the train at that new time and position.

OpenStudy (anonymous):

Get the acceleration from the increase in kinetic energy, which gives us v^2 = 2 a s, with v = velocity = 35 m/s a= acceleration (meters per second squared) s = distance = 300 m. When the last car passes by, the first one will have gone s = 385 m, and the same equation gives the velocity of the train at that new time and position. So finally, answering the question, acc=(velocity as (35 meters per second )squared) /( 2 times 300 meters) acc= 2.0416666666666666666666666666667 meters per sec squared vel = square root (2 * 2.0416666666666666666666666666667 *385) equals =39.649506091921666374402503076412 meters per second at the second position those are the answers to the question, I'm too lazy to round them.

OpenStudy (roadjester):

We aren't here to do the math, just to guide people to the response...